# GraphKeeper: Graph Domain-Incremental Learning via Knowledge Disentanglement and Preservation

### v1

@@ D.2 Implementation Details of Baselines
We implement all methods under the GIL benchmark zhang2022cglb .
@@ E.1 Comparison of Methods with Unified GCN Backbone
Considering that some baselines adopt simplified GNN such as SGC wu2019simplifying and APPNP gasteiger2018predict as their backbone, they achieve stability at the cost of sacrificing model plasticity. However, constraining the backbone of the model imposes limitations on practical applications, especially when the model architecture cannot be replaced. Thus, we unity the backbone of related methods to GCN for a fair comparison. The results are summarized in Table E.3. It show that after unifying the backbone, the performance of the baselines declines to a certain extent, significantly reducing their usability. This also highlights the advantage of our approach, which imposes no restrictions on model architecture, achieving a dual benefit of plasticity and stability.
@@ F Further Discussion on Related Work
In Sec. 4.3, we introduce the domain-aware distribution discrimination to match graphs with unobservable domains to previous domain prototypes. Recent work niu2024replay discusses a similar issue, they propose applying Laplace smoothing on original features and matching the test graph with the class prototypes through cosine similarity. Unfortunately, as shown in Figure G.3, the features across multiple domains are at risk of being excessively similar or even overlapping, which may cause the test graph to be matched with the wrong domain prototype. We guarantees a sufficiently distance between the prototypes of different domains through high-dimensional random mapping, use distance metrics in high-dimensional space as the criteria for matching, effectively preventing confusion in domain discrimination.

### v2

@@ D.2 Implementation Details of Baselines
We implement all methods under the GIL benchmark Zhang et al. (2022a).
@@ E.1 Comparison of Methods with Unified GCN Backbone
Considering that some baselines adopt simplified GNN such as SGC Wu et al. (2019) and APPNP Gasteiger et al. (2018) as their backbone, they achieve stability at the cost of sacrificing model plasticity. However, constraining the backbone of the model imposes limitations on practical applications, especially when the model architecture cannot be replaced. Thus, we unity the backbone of related methods to GCN for a fair comparison. The results are summarized in Table E.3. It show that after unifying the backbone, the performance of the baselines declines to a certain extent, significantly reducing their usability. This also highlights the advantage of our approach, which imposes no restrictions on model architecture, achieving a dual benefit of plasticity and stability.
@@ F Further Discussion on Related Work
In Sec. 4.3, we introduce the domain-aware distribution discrimination to match graphs with unobservable domains to previous domain prototypes. Recent work Niu et al. (2024a) discusses a similar issue, they propose applying Laplace smoothing on original features and matching the test graph with the class prototypes through cosine similarity. Unfortunately, as shown in Figure G.3, the features across multiple domains are at risk of being excessively similar or even overlapping, which may cause the test graph to be matched with the wrong domain prototype. We guarantees a sufficiently distance between the prototypes of different domains through high-dimensional random mapping, use distance metrics in high-dimensional space as the criteria for matching, effectively preventing confusion in domain discrimination.